One season, the average little league baseball game averaged 2 hours and 54 minutes (174 minutes) to complete. Assume the lenThe answers are in red. I completely forgot how to approach this type of question. Please go into detail!
Answer #1

solution:- le = 174 o= 13 7 =X-l 1xc190) = P(Z 190-179, =*44 190 e 190-174 13 =P(2c 1.23) - 0.8907 { from t-table? 0.8907 b)